OpenStudy (anonymous):

solve the equation. x^3 - 1=0. i know it means find the third roots of 1.

OpenStudy (anonymous):

the set up is supposed to look like this for ex: [32(cos(2pi) + isin(2pi)]^1/5

jimthompson5910 (jim_thompson5910):

you factor the left side using the difference of cubes rule x^3 - 1=0 x^3 - 1^3 =0 (x-1)(x^2 + x*1 + 1^2) = 0 (x - 1)(x^2 + x + 1) = 0

jimthompson5910 (jim_thompson5910):

then you use the zero product property to get to x-1 = 0 or x^2 + x + 1 = 0

OpenStudy (anonymous):

we didn't learn it that way for this type of problem. its power and roots of complex numbers

jimthompson5910 (jim_thompson5910):

another method x^3 - 1 = 0 x^3 = 1 x^3 = 1 + 0i 1 + 0i is in rectangular or cartesian form what is 1+0i in trignometric form?

OpenStudy (anonymous):

wouldn't the answer be 1 then?

jimthompson5910 (jim_thompson5910):

in other words, are you able to convert 1+0i to the form r*[ cos(theta) + sin(theta) ] ?

OpenStudy (anonymous):

i think so cuz i did that

jimthompson5910 (jim_thompson5910):

what did you get

OpenStudy (anonymous):

3

jimthompson5910 (jim_thompson5910):



jimthompson5910 (jim_thompson5910):

1+0i is at the point (1,0) 

jimthompson5910 (jim_thompson5910):

how far is that point away from the origin? what is the angle?

OpenStudy (anonymous):

1 point away n 0 is the angle

OpenStudy (anonymous):

is theta 0

jimthompson5910 (jim_thompson5910):

so r = 1 and theta = 0

OpenStudy (anonymous):

i get that

jimthompson5910 (jim_thompson5910):

1+0i = 1*[cos(0) + sin(0)]

jimthompson5910 (jim_thompson5910):

you'll then use de moivre's theorem

OpenStudy (anonymous):

but the root is 3 so doesn't root three go at the end?

OpenStudy (anonymous):

oh wait nevermind, no you dont

OpenStudy (michele_laino):

please you ahve to use complex numbers, so please write 1 as below: \[1 = \cos \left( {2\pi + 2k\pi } \right) + i\sin \left( {2\pi + 2k\pi } \right)\]

OpenStudy (michele_laino):

oops... you have...

OpenStudy (michele_laino):

hint: x^3-1=0, becomes: \[x=\sqrt[3]{1}\], or: \[\begin{gathered} x = \cos \left( {\frac{{2\pi + 2k\pi }}{3}} \right) + i\sin \left( {\frac{{2\pi + 2k\pi }}{3}} \right) \hfill \\ k = 0,1,2 \hfill \\ \end{gathered} \]

OpenStudy (michele_laino):

now, please substitute k=0,1,2 into the above formula and you will get your three roots, each for each value of k

OpenStudy (anonymous):

ok thnx

OpenStudy (michele_laino):

thanks!